OpenStudy (anonymous):

what is f=x{98X-3x}4y is?

hero (hero):

What do you have to do with this?

hero (hero):

I'm confused

OpenStudy (anonymous):

Im confused too ^_

hero (hero):

What are the directions?

hero (hero):

Something about your set up doesn't appear to be right

OpenStudy (anonymous):

im sorry my professor doesn't give any direction... :(

OpenStudy (anonymous):

The RHS reduces to:\[380 x^2 y \]

OpenStudy (anonymous):

what is that can u explain that to me?

hero (hero):

What? steps please?

OpenStudy (anonymous):

?

hero (hero):

Wait...is the original problem f = x(98x - 3x)4y ?

hero (hero):

Because if it is, then f = x(95x)4y = (95x^2)4y = 380x^2y

hero (hero):

I was thrown off by the 98X - 3x

OpenStudy (anonymous):

is that the answer?

hero (hero):

apparently. I was just clarifying

hero (hero):

at least trying to

OpenStudy (anonymous):

The problem statement capital X was changed to lower case x.\[x(98x-3x)4y \]\[\left(95 x^2\right)4y \]\[380 x^2 y \]